Gold has a face-centered cubic arrangement with a unit cell edge length of 4.08 Å . How many moles of gold fit in a gold nanoparticle sheet with a length of 82.8 nm , a width of 28.6 nm , and a thickness of 9.44 nm ?
Answer #1

Answer

2.19 ×10-18 mol

Explanation

Unit cell edge length = 4.08A​​​​​​= 0.408nm

volume of Unit cell = (0.408nm)3 =0.067917nm3

Volume of nonoparticle sheet = 82.8nm × 28.6nm × 9.44nm = 22355nm3

Number of unit cells in nanoparticle sheets = 22355nm3/0.067917nm3 = 329152 unit cells

Number of atoms in a face centred unit cell = 4

Numer of gold atoms in the nano sheet = 329152 × 4 = 1316608atoms

1mole = 6.022×1023 atoms

Number of moles of Gold fit to the nano sheet = (1mol/6.022×1023atoms) × 1316608atoms = 2.19× 10-18mol